# SPECIAL RELATIVITY

### v9

@@ 5.8 A Single Moving Emitter in a Noninertial Frame of Reference
Let us now consider the most general case when the system $S_{n}$ in the frame $S$ accelerated from the rest up to velocity $v$ along the $x$ axis and emitter in the accelerated frame $S_{n}$ is accelerated from rest up to velocity $u$ along the $x_{n}$ axis. Suppose that an observer in the accelerated frame $S_{n}$ performs an aberration measurement. We can describe radiation of the moving with velocity $u$ in the accelerated frame emitter in therms of the accelerated observer if we describe the electrodynamics by using metric $ds^{2}=c^{2}[1-(v-u)^{2}/c^{2}]dt_{n}^{2}-2(v-u)dx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$. At close look at the physics of this subject shows that the aberration increment is connected with the problem parameters by the relation $\theta_{a}=-v/c+u/c$. The point is that the crossed term in metric Eq.(15), which generates the anisotropy in the accelerated frame, depends only on the velocity $v$ of the accelerated frame relative to the initial inertial frame.
Figure 18: Aberration of light in an accelerated frame of reference $S_{n}$. Emitter is at rest with respect to the fixed stars. The aberration of light might be calculated from classical theory, in which the light beam may be treated, say, as a corpuscular (raindrops) beam. On the basis of classical theory one gets a light aberration angle that is only one-half as big as that predicted by special relativity.
Figure 19: The deflection of starlight by the field of the sun. The light deflection might be calculated from classical theory, in which the light beam may be treated, say, as a comet approaches with the velocity of light. On the basis of classical theory one gets a deflection of light that is only one-half as big as that predicted by general relativity.
5.9 A Useful Analogy
We have presented theoretical evidence of an aberration increment $2v/c$ concerning the single emitter problem in a non-inertial frame of reference. First, we notice that there is an extra factor 2. On the basis of classical theory one gets aberration increment that is only one-half as big as that predicted by special relativity, Fig. 18. According to classical approach, there is no principle difference between the aberration of light and the aberration of raindrops.
Now let us consider a completely different situation which, in the classical case, gives deflection of light that is only one-half as big as that predicted by special relativity. There is a methodological analogy between the aberration of light effect and the deflection of light in general relativity. A ray of light which, coming from fixed star, passes close by the sun will thus be attracted to it and will describe a somewhat concave orbit with respect to the sun. This deflection might be calculated from Newton’s theory, in which the ray of light may be treated, say, as an asteroid which approaches with the velocity of light. We then get a formula similar to that of Einstein, but giving only half the value of the deflection, Fig. 19. Errors inherent in the classical method applied to the computation of the deflection of light are due to the use of the velocity of light simply as a classical velocity. According to classical approach, there is no principal difference between the deflection of light and the deflection of asteroid. A methodological analogy with the calculation of the aberration increment from classical (raindrops) theory emerges. Does this analogy have any physical meaning? To see whether it does, we should compare the motion of relativistic particles in the gravitation field with the aberration of particle effect. We shall do this in the Section 7.5.
@@ 5.9 A Useful Analogy
Figure 20: Aberration of light in an accelerated frame of reference. Point source is at rest in the accelerated frame. Wavefront orientation presented in the absolute time coordinatization for the screen.
@@
Figure 21: Aberration of light in an accelerated frame of reference. Point source is at rest with respect to the fixed stars.
5.10 A Point Source in a Non-inertial Frame of Reference
@@
Now let us return to observations of a non-inertial observer. The peculiarity of point-like sources is that radiation emitted at one instant form a sphere around the source and the measuring instrument always influences the measured radiation, Fig. 20-Fig. 21 . The aberration of point source is considered independent of the source speed and to have just a local origin exclusively based on the observer (with his measuring instruments) speed relative to the fixed stars. The crossed term in metric Eq.(15) generates anisotropy in the accelerated frame that is responsible for the change of transmitted radiation direction (2). The point source follows the same pattern as fixed stars under the same aberration angle i.e. its apparent position changes with an angular displacement common to fixed stars.
@@
In the framework of the conventional theory of the aberration of light, there is an outstanding puzzle concerning the stellar aberration. There are double star systems the components of which change their velocity on a time scale ranging from days to years. The components of such binary systems at some times can have velocities relative to the earth very different from one another; nevertheless it is well known that these components exhibits always the same aberration angle. Rotating binary systems follow the same pattern as all fixed stars and are observed within a period of a year under the same universal aberration angle, i.e. their apparent position changes with an annual period common to all distant stars. This argument suggests that results of the astronomical observations confirm our prediction for aberration of light from a point source in an non-inertial frame. In the next chapter we discuss this experimental test in more detail.
In regard to light aberration one should differentiate between that from the point source and that from the laser source. Let us analyze the aberration of light transmitted from the laser resonator which is at rest in the accelerated frame. A close examination of of all experiments inside the accelerated frame shows that all optical phenomena in an optical resonator appeared to be independent of the acceleration relative to the fixed stars. It should be note that there is no influence of the difference between Langevin and Minkowski metrics on the parameter of a laser source. It is not hard to understand this result. It is clear that the crossed term in the Langevin metric, which generates aberration, cancels during the (round-trip) evolution of the radiation in the laser resonator. We will analyze the aberration of light transmitted from the laser resonator in great detail in the Section 8.6
@@ 5.11 Tilted Incoming Plane Wave
We go back now to the situation shown in Fig.20 - 21. A point source in the far zone produces in front of a hole aperture effectively a plane wave. In this chapter we considered only the case in which the point source on the optical axis (which is parallel to the $z_{n}$ axis) generates normally incident plane wave. It is obvious that the electrodynamics equations must be identical for all electromagnetic waves. In other words, the dispersion equation in the accelerated frame should be applied and kept in a consistent way for both incoming and scattered waves.
@@
A dispersion equation in the case of Maxwell’s electrodynamics is reduced to $k_{x}^{2}+k_{y}^{2}+k_{z}^{2}-\omega^{2}=0$. The simplicity of considered case is that even after the Galilean transformation along the $x$-axis the dispersion equation in the accelerated frame will have the same (diagonal) form $k_{z}^{2}-\omega_{i}^{2}=0$ for the normal incident wave. Now let us see what happens with incoming wave which has rate of change of phase with respect to $x_{n}$. In other words, there is the angle $\alpha$ between the wave number vector $\vec{k}$ and the $z_{n}$-axis. For the case when the angle is small we can write $\alpha\simeq k_{x}/k_{z}$. Our problem now is to work out the parameters of the incoming wave in terms of an offset $x_{o}$ of the point source with respect to the optical axis. We can write the angle of wave tilt as $\alpha\simeq x_{o}/z_{s}$, where $z_{s}$ is the distance from the source to the screen. Let us now analyze the aberration of light radiated by a point source which is at rest with respect to the accelerated frame. The wavenumber vector of the incoming plane wave is fixed by initial conditions. Let us assume that the wavefront of the emitted light beam is tilted by an angle $\alpha$ before the acceleration. Then we have $k_{x}=\alpha k$, $k_{z}=\sqrt{\omega^{2}_{i}/c^{2}-k_{x}^{2}}$. Acceleration has an effect on the field equations. With a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ equation Eq.(16) becomes: $k_{x}^{2}-2vk_{x}\omega_{i}/c^{2}+k_{z}^{2}-(\omega_{i}-vk_{x})^{2}/c^{2}=0$ in the first order approximation with respect to a small parameter $v/c$. The screen containing a hole is a kind of diffraction grating which breaks up the incident plane wave into a number of diffracted plane waves. So the wave-number vector of the transmitted plane wave along the $x_{n}$ direction is $k_{x}+K_{\perp}$ where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated transmittance. From the dispersion equation, we find the requirement that the wavenumber $K_{\perp}$ and the frequency change $\Delta\omega$ are related by $\omega_{i}\Delta\omega=-\omega_{i}K_{\perp}v+c^{2}k_{x}K_{\perp}$. The last equations state that transmitted light beam moves along the $x_{n}$ direction with group velocity $d\omega/dk_{x}=-v+c^{2}k_{x}/\omega_{i}$. Then the apparent angular position of the point source is connected with the problem parameters by the relation $\theta=-v/c+\alpha$.
Let us now look again at our dispersion equation. At first site the Doppler effect problem in the accelerated frame demonstrated the essential asymmetry between the accelerated and inertial observers. The question arises whether it is possible to determine experimentally the state of motion of the accelerated frame $S_{n}$ in the initial inertial frame $S$ by means of the Doppler shift $\Delta\omega=vk_{x}$ for accelerated source. In the Chapter 10 we demonstrate that there exist a specific feature of the Doppler shift measurements which prevents us to determine the velocity of the accelerated frame with respect to the initial inertial frame.
@@ 5.12 Absence of Reciprocity in the Aberration of Light Theory
In the present section we shall continue our discussion of the aberration of light radiated by a single moving emitter. Imaging that there are two identical emitters. Let us consider the case when the first emitter is at rest in an inertial frame and the second emitter is accelerated from rest up to velocity $v$ along the $x$ axis. Suppose that an observer, which is at rest with respect to the inertial frame of reference performs the direction of the energy transport measurement. Now we must be careful about initial phasing of these emitters. As example, we consider the case in which initially the velocity component of the light beam along the $x$-axis is equal to zero. Then how does the light beam from the moving emitter looks? The inertial observer would find that angular displacement is equal to $\theta_{a}=v/c$. That is the radiation appears as shown in Fig. 22. This is example what called the phenomenon of aberration of light and it is well known.
Now let us return to observation of an accelerated observer. When the accelerated system starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed. In this synchronization, the accelerated observer describes the light beam from an emitter at rest using usual Maxwell’s equations. One may ask: ”Where is the information about the observer acceleration recorded in the case of the Lorentz clock synchronization?” When the accelerated observer performed the standard procedure of clock synchronization, the time shift has the effect of rotation the plane of simultaneity, that is source radiation wavefront, on the angle $-v/c$. It is easy to show that information about the observer acceleration is recorded in the phase front orientation (with respect to the coordinate axes of the accelerated frame) of the light beam radiated from the source at rest in the accelerated frame.
Figure 22: The aberration of light from stationary and moving sources. According to the asymmetry between the initial inertial and accelerated frames, there is a remarkable prediction on the theory of the aberration of light. Namely, if the emitter is at rest relative to the fixed stars and the observer with measuring devices started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the emitter seen in the accelerated frame would jump by angle $\theta_{a}=-2v/c$. This situation is not symmetrical with respect to the change of the reference frames. If the observer is at rest relative to the fixed stars and the emitter started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the emitter seen in the inertial frame would jump by angle $\theta_{a}=v/c$.
@@
Figure 23: Reciprocity in a theory of the aberration of light. One must use the same source phasing to demonstrate the reciprocity. After the redirection procedure the inertial observer would find that angular displacement is $\theta_{a}=2v/c$.
@@
Let us describe what happens when accelerated observer performed the redirection of the accelerated emitter. We consider the case in which finally the (group) velocity component of the light beam along the $x_{n}$-axis is equal to zero. After this redirection procedure the inertial observer would find that angular displacement is $\theta_{a}=2v/c$. Suppose that the accelerated observer also performs an aberration measurement. Fig. 23 shows that the aberration increment is connected with the problem parameters by the relation $\theta_{a}=-2v/c$.
The situation can be described quite naturally in the following way. In this case we define reference directions using light beams. We demonstrated that symmetry is a correct concept in the measurements of the angular displacement between the inertial $z$ and accelerated $z_{n}$ reference directions as it must be.
@@
Figure 24: Procedure of the accelerated source redirection. The information about the source acceleration is recorded in the point source offset with respect to the optical axis (which is parallel to the $z_{n}$ axis). The wavefront orientation has no exact objective meaning. Because of this, the explanation of the effect of aberration (for sketch simplification) is based on the use of the absolute time (left) and Lorentz (right) coordinatization to describe how the direction of a light beam depends on the source offset.
@@
Let us return to the redirection procedure of the accelerated source. Now the question is, where is the information about the source acceleration recorded in the case of redirecting light of the accelerated emitter. In order to understand the redirection of the light source in an accelerated system, we have to watch the machinery of the source and see what happens when the redirection takes place. Fig. 13 shows a drawing of the ”plane wave” emitter type. Electrodynamics offers a procedure of redirection based on the offset of the point source with respect to the optical axis (which is parallel to the $z_{n}$ axis), Fig. 24. The information about the source acceleration is recorded in the point source offset with respect to the optical axis.
It is generally believed that the special relativity is reciprocal theory. The aberration of light problem in the accelerated frame demonstrated the essential asymmetry between the accelerated and inertial observers. In fact, without looking at anything external to the accelerated frame, one could determine the speed of the accelerated frame with respect to the initial inertial frame by means of aberration of point source measurements. No one has ever done such experiment, but we know what would happen from the astronomical observations. The motion of the stars with respect to the earth is never followed by any aberration. The aberration shift exist even in the case when star moves with the same velocity as the earth. There is an optical similarity between aberration of light from distant star (which is moving with the earth velocity) and from the earth-based point source. We will discuss this subject further in the Chapter 8.
One finds many books which say that we cannot distinguish by any experiment which observer remains ”at rest” during the acceleration, because according to the principle of relativity only relative motion has any physical meaning. It follows that any observed asymmetry would lead to a contradiction with the principle of relativity. This argument is wrong. We discuss the apparent conflict between aberration of light and the principle of relativity in the Chapter 12. It is demonstrated that there is no conflict between the fundamental structure of special relativity on the one hand, and the aberration of light phenomena. Principle of special relativity is irrelevancy of velocity with regard to physical laws, not with regard to anything.
@@ 5.13 Two Metrics
There is an important theorem in the special relativity which is this: as a result of the Galilean boost, we obtain metric of the accelerated emitter in the initial inertial frame Eq. 11. In this coordinatization, we describe the non-accelerated light source using diagonal (Minkowski) metric. We transform coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer using passive Galilean transformation: we substitute $x\to x-vt,t\to t$ into Minkowski metric of the inertial observer to obtain Eq.11. This metric describes the electrodynamics of the moving emitter with viewpoint of the inertial observer measurements.
@@
The Langevin metric describes the electrodynamics of the accelerated emitter with viewpoint of the accelerated observer measurements as viewing this of the inertial observer. We transform coordinates of the inertial observer using the inverse Galilean transformation: we substitute $x\to x_{n}+vt,t\to t_{n}$ into Minkowski metric of the inertial observer to obtain Langevin metric Eq.15. Now we see that the Langevin metric Eq.15 of the accelerated emitter in the accelerated frame is the same as metric Eq.11 of the accelerated emitter in the inertial frame. In fact the difference is only that the sign of velocity $v$. These two metrics are actually two sides of the same coin. The point is that we describe the electrodynamics of the same accelerated (with respect to the fixed stars) emitter from the point of view of both an inertial observer and an accelerated observer. One must take into account that metric tensor must be continuous quantity. This problem in special relativity can be adequately treated only by an approach uses absolute time coordinatization. For this purpose, it is sufficient to relate the coordinates and time by the Galilean boosts.
@@
The commonly accepted approach to special relativity does not account for this theorem. The metric Eq.11 is represented here for the first time. It is difficult to see why this should be, since Langevin metric (which is found by matching the accelerated frame and inertial frame metric tensors by the Galilean boost) is widely used in the applications to the optical phenomena in a rotation frame.
@@
In the problem discussed we have still a puzzle. We create a so-called comoving coordinate system in the initial inertial frame, and then discussing radiation from the moving source in terms of the new (comoving) coordinate labels. After the passive transformation, the coordinates of the source transform as $x^{\prime}=x-vt,t^{\prime}=t$. According to the equivalence of the active and passive pictures within a single inertial frame, in our case of interest the Maxwell’s equations are always valid in the comoving coordinatization. The electric field $\vec{E}^{\prime}$ of an electromagnetic wave satisfies the equation $\Box^{\prime 2}\vec{E}^{\prime}=\nabla^{\prime 2}\vec{E}^{\prime}-\partial^{2}\vec{E}^{\prime}/\partial(ct^{\prime})^{2}=0$. In other words, Minkowski metric $ds^{2}=c^{2}(dt^{\prime})^{2}-(dx^{\prime})^{2}-(dy^{\prime})^{2}-(dz^{\prime})^{2}$ describes the electrodynamics of the accelerated light source in comoving coordinatization. We will then ask how it is possible to take Minkowski metric for comoving coordinatization in the inertial frame and the Langevin metric in the accelerated frame. This approach needs an explanation.
@@
The difference between measurements made in both reference frames is very interesting. First, we must now discuss a certain feature of the correspondence between passive and active transformations. It should be note that passive Galilean boost within a single inertial frame, $x^{\prime}=x-vt,t^{\prime}=t$, is simply another parametrization of the observations of the inertial observer. The Eq.11 describes the inertial observer measurements. According to active transformations, the motion of fixed stars with respect to the inertial observe and his devices not changes and there is equivalence of the active and passive Galilean boost. In contrast, the Langevin metric describes the accelerated observer measurements. According to the inverse Galilean boost, the motion of fixed stars with respect to the space-time grid $(t_{n},x_{n},y_{n},z_{n})$ of the accelerated observer changes.
@@
On the other hand, the standard textbook presentation of the relativistic transformation is based on the hidden assumption that $x_{n}y_{n}z_{n}$-axes of the accelerated observer and $xyz$-axes of the inertial observer are parallel. In other words, it is assumed that observers have common 3-space. It is generally believed that simultaneous acceleration of the (rigid) reference space grid has direct physical meaning. This is misconception. The commonly accepted derivation of the aberration of light effect does not account for the conventionality of distant simultaneity. There is an uncertainty of the relative position along the $z_{n}$-axis in the $x$-direction of amount $z_{n}v/c$ due to the uncertainty in the moments of acceleration. In fact in this case we have deal with space-like separated events. The $z_{n}$ axis orientation has no exact objective meaning, because the relativity of simultaneity takes place. The statement that the $z_{n}$-axis orientation with respect to the $x$-axis of the inertial observer has objective meaning to within a certain accuracy can be visualized by the picture of $z_{n}$ axis in the vertical orientation with approximate angle extension (blurring) given by $\Delta\theta=v/c$. This inclination angle between the inertial $z$ and accelerated $z_{n}$ reference directions has no exact objective meaning at all since, due to finiteness of the speed of light, we cannot specify any experimental method by which this displacement could be ascertained. We shall discuss this subject further in the Chapter 7.

### v10

@@ 5.8 A Single Moving Emitter in an Accelerated Frame of Reference
Let us now consider the most general case when the system $S_{n}$ in the frame $S$ accelerated from the rest up to velocity $v$ along the $x$ axis and emitter in the accelerated frame $S_{n}$ is accelerated from rest up to velocity $u$ along the $x_{n}$ axis. Suppose that an observer in the accelerated frame $S_{n}$ performs an aberration measurement. We can describe radiation of the moving with velocity $u$ in the accelerated frame emitter in terms of the accelerated observer if we describe the electrodynamics by using metric $ds^{2}=c^{2}[1-(v-u)^{2}/c^{2}]dt_{n}^{2}-2(v-u)dx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$. A close look at the physics of this subject shows that the aberration increment is connected with the problem parameters by the relation $\theta_{a}=-v/c+u/c$. The point is that the crossed term in metric Eq.(15), which generates the anisotropy in the accelerated frame, depends only on the velocity $v$ of the accelerated frame relative to the initial inertial frame.
@@
Figure 18: Aberration of light in an accelerated frame of reference. A point source is at rest in the accelerated frame. Wavefront orientation is presented in the absolute time coordinatization for the screen.
@@
Figure 19: Aberration of light in an accelerated frame of reference. A point source is at rest with respect to the fixed stars.
5.9 A Point Source in an Accelerated Frame of Reference
@@
Now let us return to observations of an accelerated observer. The peculiarity of point-like sources is that radiation emitted at one instant forms a sphere around the source and the measuring instrument always influences the measured radiation, Fig. 18 - Fig. 19. The aberration of a point source is considered independent of the source speed and to have just a local origin exclusively based on the observer’s speed relative to the fixed stars. The crossed term in metric Eq.(15) generates anisotropy in the accelerated frame that is responsible for the change of transmitted radiation direction. The point source follows the same pattern as fixed stars under the same aberration angle i.e. its apparent position changes with an angular displacement common to fixed stars.
@@
In the framework of the conventional theory of the aberration of light, there is an outstanding puzzle concerning stellar aberration. There are double star systems the components of which change their velocity on a time scale ranging from days to years. The components of such binary systems at some times can have velocities relative to the earth very different from one another; nevertheless, it is well known that these components always exhibit the same aberration angle. Rotating binary systems follow the same pattern as all fixed stars and are observed within a year under the same universal aberration angle, i.e. their apparent position changes with an annual period common to all distant stars. This argument suggests that the results of the astronomical observations confirm our prediction for the aberration of light from a point source in a rotating frame. In the next section, we discuss this experimental test in more detail.
@@ 5.10 Tilted Incoming Plane Wave
We go back now to the situation shown in Fig.18 - Fig. 19. A point source in the far zone produces in front of a hole aperture effectively a plane wave. In this section, we considered only the case in which the point source on the optical axis (which is parallel to the $z_{n}$ axis) generates a normally incident plane wave. The electrodynamics equations must be identical for all electromagnetic waves. In other words, the dispersion equation in the accelerated frame should be applied and kept consistently for both incoming and scattered waves.
@@
A dispersion equation in the case of Maxwell’s electrodynamics is reduced to $k_{x}^{2}+k_{y}^{2}+k_{z}^{2}-\omega^{2}=0$. The simplicity of the considered case is that even after the Galilean transformation along the $x$-axis the dispersion equation in the accelerated frame will have the same (diagonal) form $k_{z}^{2}-\omega_{i}^{2}=0$ for the normal incident wave. Now let us see what happens with an incoming wave which has a rate of change of phase with respect to $x_{n}$. In other words, there is the angle $\alpha$ between the wave number vector $\vec{k}$ and the $z_{n}$-axis. For the case when the angle is small we can write $\alpha\simeq k_{x}/k_{z}$. Our problem now is to work out the parameters of the incoming wave in terms of an offset $x_{o}$ of the point source with respect to the optical axis. We can write the angle of wave tilt as $\alpha\simeq x_{o}/z_{s}$, where $z_{s}$ is the distance from the source to the screen. Let us now analyze the aberration of light radiated by a point source which is at rest with respect to the accelerated frame. The wavenumber vector of the incoming plane wave is fixed by initial conditions. Let us assume that the wavefront of the emitted light beam is tilted by an angle $\alpha$ before the acceleration. Then we have $k_{x}=\alpha k=\alpha\omega_{i}/c$, $k_{z}=\sqrt{\omega_{i}^{2}/c^{2}-k_{x}^{2}}$. Acceleration affects the field equations. With a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ equation Eq.(17) becomes: $(1-v^{2}/c^{2})k_{x}^{2}-2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$. The screen containing a hole is a kind of diffraction grating that breaks up the incident plane wave into a number of diffracted plane waves. So the wave-number vector of the transmitted plane wave along the $x_{n}$ direction is $k_{x}=\alpha k+K_{\perp}$ where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated transmittance. In keeping with our approximation, we should leave out the terms $v^{2}/c^{2}$, $\alpha^{2}$, and $\alpha v/c$. Our dispersion equation now looks like $\omega_{i}\Delta\omega=-\omega_{i}K_{\perp}v+c^{2}\alpha kK_{\perp}$. The last equations state that the transmitted light beam moves along the $x_{n}$ direction with group velocity $d\omega/dk_{x}=-v+c^{2}\alpha k/\omega_{i}$. Then the apparent angular position of the point source is connected with the problem parameters by the relation $\theta=-v/c+\alpha$ [^26] A similar situation would occur if we consider the case of a screen, in the inertial frame, moving with velocity $v$ along the $x$-axis. In Chapter 4, we have taken a simple example where a plane wave falls normally on the screen. But of course, we can generalize it. Let us assume that the wavefront of the incoming plane wave is tilted by an angle $\alpha$. According to the presented approach, there is a deviation of the energy transport for light transmitted through a hole in the moving screen. The aberration increment is connected with problem parameters by the relation $\theta=v/c+\alpha$..
@@ 5.11 Absence of Reciprocity in the Aberration of Light Theory
In the present section, we shall continue our discussion of the aberration of light radiated by a single moving emitter. Imagine that there are two identical emitters. Let us consider the case when the first emitter is at rest in an inertial frame and the second emitter is accelerated from rest up to velocity $v$ along the $x$ axis. Suppose that an observer, which is at rest with respect to the inertial frame of reference performs the direction of the energy transport measurement. Now we must be careful about the initial phasing of these emitters. For example, we consider the case in which initially the velocity component of the light beam along the $x$-axis is equal to zero. Then how does the light beam from the moving emitter look? The inertial observer would find that angular displacement is equal to $\theta_{a}=v/c$. That is the radiation appears as shown in Fig. 20. This is an example of what is called the phenomenon of aberration of light and it is well known.
Figure 20: The aberration of light from stationary and moving sources.
@@
Figure 21: Aberration of light when the accelerated observer redirected the accelerated emitter. After the redirection procedure, the inertial observer would find that angular displacement is $\theta_{a}=2v/c$.
@@
Let us describe what happens when the accelerated observer performed the redirection of the accelerated emitter. We consider the case in which finally the (group) velocity component of the light beam along the $x_{n}$-axis is equal to zero. After this redirection procedure, the inertial observer would find that angular displacement is $\theta_{a}=2v/c$. Suppose that the accelerated observer also performs an aberration measurement. Fig. 21 shows that the aberration increment is connected with the problem parameters by the relation $\theta_{a}=-2v/c$.
@@
Figure 22: The procedure of the accelerated source redirection. The information about the source acceleration is recorded in the point source offset with respect to the optical axis (which is parallel to the $z_{n}$ axis). The wavefront orientation has no exact objective meaning. Because of this, the explanation of the effect of aberration (for sketch simplification) is based on the use of absolute time (left) and Lorentz (right) coordinatization to describe how the direction of a light beam depends on the source offset.
@@
Let us return to the redirection procedure of the accelerated source. Now the question is, where is the information about the source acceleration recorded in the case of redirecting light of the accelerated emitter? To understand the redirection of the light source in an accelerated system, we have to watch the machinery of the source and see what happens when the redirection takes place. Fig.13 shows a drawing of the ”plane wave” emitter type. Electrodynamics offers a procedure of redirection based on the offset of the point source with respect to the optical axis (which is parallel to the $z_{n}$ axis), Fig.22. The information about the source acceleration is recorded in the point source offset with respect to the optical axis.
@@ 5.12 Two Metrics
There is an important theorem in special relativity which is this: as a result of the Galilean boost, we obtain a metric of the accelerated emitter in the initial inertial frame Eq. 11. In this coordinatization, we describe the non-accelerated light source using the diagonal (Minkowski) metric. We transform coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer using the equivalence of the active and passive pictures: we substitute $x\to x-vt,t\to t$ into Minkowski metric of the inertial observer to obtain Eq.11. This metric describes the electrodynamics of the moving emitter with the viewpoint of the inertial observer measurements.
@@
The Langevin metric describes the electrodynamics of the accelerated emitter with the viewpoint of the accelerated observer measurements as viewing this of the inertial observer. One must take into account that the metric tensor must be a continuous quantity. This problem in special relativity can be adequately treated only by an approach that uses absolute time coordinatization. For this purpose, it is sufficient to relate the coordinates and time to the Galilean boosts. We transform the coordinates of the inertial observer using the inverse Galilean transformation: we substitute $x\to x_{n}+vt,t\to t_{n}$ into Minkowski metric of the inertial observer to obtain Langevin metric Eq.15. Now we see that the Langevin metric Eq.15 of the accelerated emitter in the accelerated frame is the same as metric Eq.11 of the accelerated emitter in the inertial frame. The difference is only that the sign of velocity $v$. These two metrics are two sides of the same coin. The point is that we describe the electrodynamics of the same accelerated (with respect to the fixed stars) emitter from the point of view of both an inertial observer and an accelerated observer.
@@
The commonly accepted approach to special relativity does not account for this theorem. The metric Eq.11 is represented here for the first time. It is difficult to see why this should be since the Langevin metric (which is found by matching the accelerated frame and inertial frame metric tensors by the Galilean boost) is widely used in applications to the optical phenomena in a rotating frame.
@@
In the problem discussed we have still a puzzle. We create a so-called comoving coordinate system in the initial inertial frame and then discuss radiation from the moving source in terms of the new (comoving) coordinate labels. After the passive transformation, the source coordinates transform as $x^{\prime}=x-vt,t^{\prime}=t$. According to the equivalence of the active and passive pictures within a single inertial frame, in our case of interest, Maxwell’s equations are always valid in the comoving coordinatization. The electric field $\vec{E}^{\prime}$ of an electromagnetic wave satisfies the equation $\Box^{\prime 2}\vec{E}^{\prime}=\nabla^{\prime 2}\vec{E}^{\prime}-\partial^{2}\vec{E}^{\prime}/\partial(ct^{\prime})^{2}=0$. In other words, Minkowski metric $ds^{2}=c^{2}(dt^{\prime})^{2}-(dx^{\prime})^{2}-(dy^{\prime})^{2}-(dz^{\prime})^{2}$ describes the electrodynamics of the accelerated light source in comoving coordinatization. We will then ask how it is possible to take the Minkowski metric for comoving coordinatization in the inertial frame and the Langevin metric in the accelerated frame. This approach needs an explanation.
@@
The difference between measurements made in both reference frames is very interesting. First, we must discuss a certain feature of the correspondence between passive and active transformations. It should be noted that passive Galilean boost within a single inertial frame, $x^{\prime}=x-vt,t^{\prime}=t$, is simply another parametrization of the observations of the inertial observer. The Eq.11 describes the inertial observer measurements. According to active transformations within a single inertial frame, the motion of fixed stars with respect to the inertial observer and his devices does not change and there is the equivalence of the active and passive Galilean boost. In contrast, the Langevin metric describes the accelerated observer measurements. According to the inverse Galilean boost, the motion of fixed stars with respect to the space-time grid $(t_{n},x_{n},y_{n},z_{n})$ of the accelerated observer changes.
@@
On the other hand, the standard textbook presentation of the relativistic transformation is based on the hidden assumption that $x_{n}y_{n}z_{n}$-axes of the accelerated observer and $xyz$-axes of the inertial observer are parallel. In other words, it is assumed that observers have a common 3-space. It is generally believed that simultaneous acceleration of the (rigid) reference space grid has direct physical meaning. This is a misconception. The commonly accepted derivation of the aberration of light effect does not account for the conventionality of distant simultaneity. There is an uncertainty of the relative position along the $z_{n}$-axis in the $x$-direction of amount $z_{n}v/c$ due to the uncertainty in the moments of acceleration. In this case we have to deal with space-like separated events. The $z_{n}$ axis orientation has no exact objective meaning because the relativity of simultaneity takes place. The statement that the $z_{n}$-axis orientation with respect to the $z$-axis of the inertial observer has objective meaning to within a certain accuracy can be visualized by the picture of $z_{n}$ axis in the vertical orientation with approximate angle extension (blurring) given by $\Delta\theta=v/c$. This inclination angle between the inertial $z$ and accelerated $z_{n}$ reference directions has no exact objective meaning at all since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this displacement could be ascertained. We shall discuss this subject further in Chapter 7.